# Mapping global offshore wind resource: wake losses, optimisation potential and climate effects

### v1

@@ 2.3 Mapping layout optimisation potential
The distribution of wind directions at a given location influences the optimal array layout. A highly focused wind direction distribution presents the possibility of mitigating wake losses by placing downstream turbines outside of the wake of upstream turbines. For uniformly distributed wind directions, however, we expect to find significantly less potential for optimisation.
We first investigate the variation in wind direction spread across EEZs, based on the ERA-5 reanalysis data. As a measure of wind direction spread, we use the circular standard deviation of the wind direction.
We then estimate the optimisation potential at each grid cell within the model domain. This is based on taking the ‘standard’ farm as a starting point, and optimising using the TopFarm Python package (Réthoré et al., 2014). This uses PyWake to estimate power for a given array layout, and interfaces with SciPy to perform a gradient-based optimisation of the position of each turbine, to maximise total power output. We define the optimisation potential as
where is given by equation (6) and is obtained by optimisation using TopFarm. Note that for consistency with the other results we present, is computed by averaging the wake effects over all rotations of the initial wind farm, which has a square grid layout. However, the optimal layout design is based on optimising from a single initial condition in which the farm grid layout is aligned with the mean wind direction.
Since there are approximately 300,000 grid cells within the global EEZs at the resolution we consider, the computational cost associated with optimising the array layout of the 100-turbine farm for every grid cell is prohibitive. To mitigate this, we first coarsen the grid by a factor of 4 in each dimension, for a total of 18,491 grid cells. The coarsening discards data from unselected grid cells, since spatial averaging would produce artificially smooth wind roses. We also choose to perform the optimisation on a 25-turbine farm, with a 5 by 5 grid layout, rather than the full 100-turbine farm. This reduces the computational cost sufficiently to enable this global analysis. We further perform the optimisation on the full 100-turbine problem for a subset of 288 grid cells, obtained by coarsening the grid by a factor of 32 in each dimension. For each of the 25-turbine and 100-turbine problems, we obtain the optimisation potential given by equation (9). For the grid cells which appear in both datasets, we find a strong linear relationship between the optimisation potentials for the 25-turbine and 100-turbine problems, and therefore train a linear regression model to map from one to the other. Using this model, we then estimate the 100-turbine optimisation potential on the finer grid, on which we have run the 25-turbine optimisation. Further detail regarding the linear regression model is provided in A.
@@ 2.4 Mapping future changes due to climate
We restrict our analysis of climate change effects to the influence of wind speed distribution. For each climate model, we follow the method of section 2.2 to compute the ambient and ‘waked’ farm capacity factors, for both the ‘historical’ and future periods. From these, we firstly compute the projected change in ambient resource, as
where denotes the ambient CF for the future period. We secondly compute the projected change in wake-induced losses, as
@@
The results we present within this study are based on the mean changes across the ensemble of 10 climate model projections. In order to identify robust changes, we isolate grid cells for which the sign of the projected changes is the same across at least 8 of the 10 models.
@@ 3.1 Mapping wake effects
Figure 1(a) shows the global distribution of ambient capacity factors, for the 10 reference turbine, based on ERA-5 data from our ‘historical’ period of 1995–2014. Consistent with well-established global wind climatology (Zheng et al., 2018), winds are strongest in the mid-latitudes, and in particular we find very high ambient capacity factors for latitudes between -40 and -60. Wind resources are smallest in the tropics. This result constitutes the most basic form of wind resource mapping, where we are interested only in the mean ambient wind power.
@@
Figure 1(b) shows the global distribution of the capacity factor lost to intra-farm wakes for the 10 by 10 reference array averaged over 36 orientations, given by equation (7). Figure 1(c) shows the % losses given by equation (8). Comparing with figure 1(a), in general we find that regions of high (low) ambient resource coincide with % losses of low (high) magnitude. Since wake losses are negative, this amounts to a positive correlation between ambient CF and % losses.
@@
Investigating this relationship further, figure 2 shows the relationship between the ambient CF and (a) CF lost to wakes, and (b) % wake losses. The solid lines show the global mean losses, as a function of ambient CF. Focusing on figure 2(a), for low ambient CFs, we find a near-linear relationship, or equivalently in figure 2(b) the % losses initially change slowly with respect to ambient CF. However, for moderate ambient CFs of around 30% or higher, we find a bifurcation in figure 2(a). For most locations, the CF lost begins to reduce in magnitude (closer to zero) with increasing ambient CF. This can be explained by the decreasing fraction of wind speeds which fall between the cut-in and rated speeds of our reference wind turbine (3 and 11 respectively), where wake effects are most significant. Most regions with higher average wind speeds produce greater ambient power, and also spend less time in this wake-sensitive interval, resulting in a reduction in absolute CF lost. However, there are some regions for which the near-linear relationship between ambient CF and CF lost continues, resulting in much greater wake losses. Looking at regions with ambient CF of, for example, 55–60%, there is a very broad spread of CF lost, and hence net CF which is extracted by our ‘standard’ farm.
@@
Figure 3 reproduces the results of figure 2, but presents results aggregated per country, with selected countries highlighted. We observe that the United Kingdom, with access to high-resource offshore regions including in the North Sea and Irish Sea, has a high mean ambient CF of 65%, and relatively modest wake-induced losses of 4.5%. The United States, Australia and China have moderate ambient CF, and also moderate losses. We include Venezuela as a fairly extreme example of high wake losses which is fairly typical of Atlantic South America; it has a strong ambient resource of 56%, but suffers more than almost any other country in terms of absolute wake-induced losses, with 5.6% CF lost, equivalent to 10.3% losses. Brazil has the lowest ambient CF of the highlighted countries, and also suffers from greater than average losses, although is not as extreme as Venezuela.
@@
Figure 4 presents the difference between the actual % losses, and the ‘global mean’ % losses as a function of ambient CF, shown by the solid line in figures 2(b) and 3(b). In general, the mid-latitudes suffer from smaller than average wake-induced losses (hence differences to the mean are positive), while the tropics suffer from larger losses (negative differences to the mean). We emphasise that these losses are relative to the global mean for the local ambient CF. We highlight particular regions of interest in figure 5. We find that Europe, including the Baltic and Mediterranean seas, possesses strong ambient resources, and favourable wake-induced losses. The US East Coast has moderate ambient resource and also has favourable losses. The US West Coast has stronger ambient resource than the East Coast. Near-shore regions exhibit favourable losses, but for regions further offshore the losses are stronger. Australia, with such a large spatial footprint, has a very broad range of ambient CFs, although some of the higher-resource regions coincide with larger-than-average wake-induced losses. India has fairly low ambient resource but mostly favourable wake losses. China has moderate ambient CFs, with slightly smaller-than-average wake effects. Atlantic South America suffers from some of the worst wake-induced losses. As identified in figure 3, Venezuela has strong ambient resource but very large wake-induced losses. Brazil similarly suffers from significant losses, as does most of the Atlantic coast of Colombia. These two regions were considered ‘markets to watch’ by GWEC (2023); these results suggest that a different approach to planning and design may be required, compared with Europe, due to the stronger wake effects experienced.
@@
Overall, these results reveal the inadequacy of any (global) resource assessment which neglects wake effects, or which assumes spatially uniform losses. Neither absolute or percentage losses are spatially uniform. Both exhibit overall trends with respect to ambient CF, but there remains significant variation even among locations with the same ambient resource. In terms of absolute losses, there is a turning point where regions with greater ambient resource actually suffer from smaller absolute wake-induced losses; this pertains mostly to middle-latitude regions. In contrast, tropical regions with moderate to high ambient resource are likely to suffer from the largest wake-induced losses, with tropical South America particularly impacted.
@@ 3.1.1 Explaining the spatial variation in wake-induced losses
To explain the observed spatial variation in wake-induced losses, we must inspect specific wind speed distributions, in comparison to the turbine power and thrust curves. Figure 6 shows the mean wind speed distributions for Europe (including the Baltic and Mediterranean seas), and Venezuela, which was identified above as a country that experiences very high losses. The ambient capacity factors for these regions are very similar (55.8% for Europe, 56.1% for Venezuela), but the CF lost to wakes differs significantly (-2.9% for Europe, -5.6% for Venezuela). This is due to their differing wind speed distributions, as shown in figure 6. Venezuela exhibits a far narrower wind speed distribution than Europe, coinciding with the most wake-sensitive part of the turbine power and thrust curves (figure 7), which is between around 3 and 11 . In Europe, a substantial part of the wind speed distribution sits above 11 , where power output is at its maximum and is insensitive to wake effects.
This influence of the wind speed distribution on wake effects can be more formally explained by considering one-dimensional momentum theory (see e.g. Kulunk (2011)). The thrust coefficient of a turbine can be related to the axial induction factor , by
Given the wind turbine thrust curve data for our selected turbine model (see figure 7), we can thus derive as a function of upstream wind speed. Neglecting wake expansion, the reduction in wind speed downstream of a wind turbine is then given by
@@
Figure 8 shows this estimate as a function of wind speed. This is consistent with our earlier statements regarding the most important wind speed interval for wake effects, between around 3 and 11 . Figure 8 also shows the actual average power lost per turbine by the ‘standard’ farm we study within this work, as a function of wind speed, as simulated using PyWake. Although corresponds to the theoretical power lost by a single turbine downstream of another, neglecting wake expansion, it is clear that it is a good indicator of wake-induced losses at a given wind speed.
To derive a theoretical estimator for the overall mean wake-induced losses, we combine this with the wind speed distribution at a given location . Specifically, we calculate
We find that this quantity is very strongly correlated with the simulated wake losses, with an coefficient of 0.98.
We note that this formulation does not depend on the total capacity or layout of the wind farm, but only on the power and thrust curves of the selected turbine. It follows that turbine models with different thrust and power curves will produce a different spatial pattern of wake-induced losses. It may further be possible to mitigate losses by careful selection of turbine types appropriate to the wind speed distribution at the site where they will be deployed. Further exploration of the wake response for different turbine types is beyond the scope of the present study but will be considered in future work.
@@ 3.2 Mapping layout optimisation potential
Figure 10 shows the global distribution of the array layout optimisation potential, which is defined as the percentage improvement in mean power which can be obtained by an optimised layout, compared with the initial square farm layout. There is a negative correlation between wind direction standard deviation and optimisation potential. That is, narrower wind speed distributions are more conducive to array layout optimisation. We find the greatest optimisation potentials in northern South America, particularly on the western (Pacific) coast.
Comparing the optimisation potential of figure 10 with the global distribution of excess wake losses of figure 4, we find that the regions (mostly equatorial) where the wake losses are particularly high also have the greatest optimisation potential. This is also shown in figure 11, which should be compared with figure 2. Mean losses at all ambient CFs are improved by optimisation. Looking at the absolute losses (figs 2(a) and 11(a)), the optimisation has the greatest benefit for ambient CFs around 50–60%. In terms of percentage losses (figs 2(b) and 11(b)), the greatest improvements are found for ambient CFs of around 10–40%.
@@ 3.3 Mapping future changes due to climate
Figure 12(a) shows the mean projected percentage change in ambient capacity factor across the 10 climate models, between the ‘historical’ climate, and 2081–2100 under the SSP585 scenario. Only grid cells with a statistically significant change are shown, where our significance criterion is that at least 8 out of the 10 climate models show the same sign of change. The result is broadly consistent with the global analysis of Ibarra-Berastegui et al. (2023), as well as with other regional studies. We observe statistically significant decreases in offshore wind resource across both Europe (consistent with (Carvalho et al., 2021)) and the US (consistent with Martinez and Iglesias (2022a)). We find a decrease in resource offshore of Northern China, which is consistent with Zhang and Li (2021), Deng et al. (2024), although we do not find statistically significant increases in south China. We find decreasing resources on the west coast of India (Basak et al., 2023), as well as most of Australia (Fournier et al., 2023). Our projections around South America are also consistent with the study of de Souza Ferreira et al. (2024).
@@
Figures 13 and 14 show the relationship between the % projected change in ambient CF, and % losses, per grid cell (figure 13), and aggregated per country (figure 14). For the aggregation by country, the significance is determined on a per-country basis, rather than by aggregation over only the significant grid cells belonging to each country. Both indicate a positive correlation between change in ambient CF, and change in % wake losses. This is consistent with the trend found in section 3.1, and in particular figure 2(b).
@@
The majority of grid cells/countries fall into the lower left quadrant of figures 13 and 14, indicating that ambient resource decreases, but that percentage wake-induced losses increase in magnitude, exacerbating the climate effect. Figure 15 indicates the spatial distribution of these grid cells. These regions include the northern Atlantic, impacting the US, Atlantic Canada, Greenland, Iceland, Ireland and the UK, as well as the US west coast, parts of Australia, India and China, southern Chile, and a number of island states in the southern Pacific. In these regions, decreases in ambient wind resource due to climate change are exacerbated by an increase in the magnitude of the percentage wake-induced losses.
@@
In contrast, countries falling into the upper right quadrant of figures 13 and 14 are projected to both increase in ambient resource and improve in wake losses; these countries include Russia and Canada (driven mostly by strong increases in resource in the Arctic), as well as other isolated regions such as Pakistan and Oman. There are no countries which fall within either the upper left or lower right quadrants for which the statistical significance threshold is met, although there are a small number of individual grid cells meeting these criteria, as shown in figure 13.
@@
Considering the globe as a whole, the direct climate impact on ambient power tends to have the greatest magnitude among the factors considered in this study, with magnitudes exceeding 25% for some regions, such as India. The magnitude of the baseline wake-induced losses (for the current climate) did not exceed 15%. Optimisation potentials reached up to 6–7%, while the indirect climate impact via wake effects has magnitudes of around 1%. However, we note that the relative size and significance of these effects varies by region. For example, there are some regions where optimisation potential exceeds the projected impact of climate change on ambient resource. This serves to further highlight that the challenges facing the offshore wind industry are highly region-dependent.

### v2

@@ 2.3 Mapping layout optimisation potential
The distribution of wind directions at a given location influences the optimal array layout. A highly focused wind direction distribution presents the possibility of mitigating wake losses by placing downstream turbines outside of the wake of upstream turbines. For uniformly distributed wind directions, however, significantly less potential for optimisation is expected.
To explore this, the variation in wind direction spread across EEZs is first investigated, based on the ERA-5 reanalysis data. The circular standard deviation of the wind direction is used as a measure of wind direction spread.
The optimisation potential at each grid cell within the model domain is then estimated by taking the ‘standard’ farm as a starting point, and optimising using the TopFarm Python package, version 2.3.2 (Réthoré et al., 2014). This uses PyWake to estimate power for a given array layout, and interfaces with SciPy to perform a gradient-based optimisation of the position of each turbine, to maximise total power output. The optimisation potential is then defined as
where is given by equation (6) and is obtained by optimisation using TopFarm. Note that for consistency with the other results presented, is computed by averaging the wake effects over all rotations of the initial wind farm, which has a square grid layout. However, the optimal layout design is based on optimising from a single initial condition in which the farm grid layout is aligned with the mean wind direction.
Since there are approximately 300,000 grid cells within the global EEZs at the resolution considered, the computational cost associated with optimising the array layout of the 100-turbine farm for every grid cell is prohibitive. To mitigate this, the grid is coarsened by a factor of 8 in each dimension, for a total of 4,632 grid cells. The coarsening discards data from unselected grid cells, since spatial averaging would produce artificially smooth wind roses. Although the spatial resolution is decreased, this level of coarsening maintains the ability to observe global trends in optimisation potential, at a manageable computational cost.
@@ 2.4 Mapping future changes due to climate
The analysis of climate change effects in this study is restricted to the influence of wind speed distribution. This neglects any effects due to changing wind direction distribution, and also neglects any potential climatic feedback from large-scale wind deployment (Wang et al., 2023a); although both of these aspects warrant further research, they lie beyond the scope of this study. For each climate model, the method of section 2.2 is used to compute the ambient and ‘waked’ farm capacity factors, for both the ‘historical’ and future periods. From these, the projected change in ambient potential is calculated as
where denotes the ambient CF for the future period and denotes the ambient CF from the historical period. The projected change in wake-induced losses is then calculated as
@@
The results presented within this study are based on the mean changes in the above quantities, across the ensemble of 10 climate model projections. Note that the ‘change’ calculated for a particular model is derived from the bias-corrected future and historical data from that model (not with the ERA-5 reanalysis). A sign test is used in order to identify statistically significant changes; the projected change in a given grid cell is statistically significant if the sign of the projected changes is the same across at least 8 of the 10 models. This corresponds to a value of 0.055. The sign test is selected since it is a robust statistical test even for small samples sizes, such as in this case where .
@@ 3.1 Mapping wake effects
Figure 2(a) shows the global distribution of ambient capacity factors, for the 10 reference turbine, based on ERA-5 data from our ‘historical’ period of 1995–2014. Consistent with well-established global wind climatology (Zheng et al., 2018), winds are strongest in the mid-latitudes, and in particular the results show very high ambient capacity factors for latitudes between -40 and -60. Wind potentials are smallest in the tropics. This result constitutes the most basic form of wind potential mapping, where focused only in the mean ambient wind power.
@@
Figure 2(b) shows the global distribution of the capacity factor lost to intra-farm wakes for the 10 by 10 reference array averaged over 36 orientations, given by equation (7). Figure 2(c) shows the % losses given by equation (8). Comparing with figure 2(a), in general, regions of high (low) ambient potential coincide with % losses of low (high) magnitude. Since wake losses are negative, this amounts to a positive correlation between ambient CF and % losses.
@@
Investigating this relationship further, figure 3 shows the relationship between the ambient CF and (a) CF lost to wakes, and (b) % wake losses. The solid lines show the global mean losses, as a function of ambient CF. Focusing on figure 3(a), for low ambient CFs, there is a near-linear relationship. However, for moderate ambient CFs of around 30% or higher, there is a bifurcation. For most locations, the CF lost begins to reduce in magnitude (closer to zero) with increasing ambient CF. This can be explained by the decreasing fraction of wind speeds which fall between the cut-in and rated speeds of our reference wind turbine (3 and 11 respectively), where wake effects are most significant. Most regions with higher average wind speeds produce greater ambient power, and also spend less time in this wake-sensitive interval, resulting in a reduction in absolute CF lost. However, there are some regions for which the near-linear relationship between ambient CF and CF lost continues, resulting in much greater wake losses. Looking at regions with ambient CF of, for example, 55–60%, there is a very broad spread of CF lost, and hence net CF which is extracted by our ‘standard’ farm.
@@
Figure 4 reproduces the results of figure 3, but presents results aggregated per country, with selected countries highlighted. The United Kingdom, with access to high-resource offshore regions including in the North Sea and Irish Sea, has a high mean ambient CF of 65%, and relatively modest % wake-induced losses of 8.3%. The United States, Australia and China have moderate ambient CF, and also moderate % losses of around 13%. Venezuela is included as a fairly extreme example of high wake losses which is fairly typical of Atlantic South America; it has a strong ambient potential of 56%, but suffers more than almost any other country in terms of absolute wake-induced losses, with 10.1% CF lost, equivalent to 18.5% losses. At 41.7%, Brazil has the lowest ambient CF of the highlighted countries, and also suffers from greater than average % losses of 19.3%.
@@
Figure 5 presents the difference between the actual % losses, and the ‘global mean’ % losses as a function of ambient CF, shown by the solid line in figures 3(b) and 4(b). In general, the mid-latitudes suffer from smaller than average wake-induced losses (hence differences to the mean are positive), while the tropics suffer from larger losses (negative differences to the mean). It should be emphasised that these losses are relative to the global mean for the local ambient CF. Selected regions of interest are highlighted in figure 6. Europe, including the Baltic and Mediterranean seas, possesses strong ambient potential, and favourable wake-induced losses. The US East Coast has moderate ambient potential and also has favourable losses. The US West Coast has stronger ambient potential than the East Coast. Near-shore regions exhibit favourable losses, but for regions further offshore the losses are stronger. Australia, with such a large spatial footprint, has a very broad range of ambient CFs, although some of the higher-resource regions coincide with larger-than-average wake-induced losses. India has fairly low ambient potential but mostly favourable wake losses. China has moderate ambient CFs, with slightly smaller-than-average wake effects. Atlantic South America suffers from some of the worst wake-induced losses. As identified in figure 4, Venezuela has strong ambient potential but very large wake-induced losses. Brazil similarly suffers from significant losses, as does most of the Atlantic coast of Colombia. These two regions were considered ‘markets to watch’ by GWEC (2023); these results suggest that a different approach to planning and design may be required, compared with Europe, due to the stronger wake effects experienced.
@@
Overall, these results reveal the inadequacy of any (global) resource assessment which neglects wake effects, or which assumes spatially uniform losses. Neither absolute or percentage losses are spatially uniform. Both exhibit overall trends with respect to ambient CF, but there remains significant variation even among locations with the same ambient wind potential. In terms of absolute losses, there is a turning point where regions with greater ambient potential actually suffer from smaller absolute wake-induced losses; this pertains mostly to middle-latitude regions. In contrast, tropical regions with moderate to high ambient potential are likely to suffer from the largest wake-induced losses, with tropical South America particularly impacted.
@@ 3.1.1 Explaining the spatial variation in wake-induced losses
To explain the observed spatial variation in wake-induced losses, it is necessary to inspect specific wind speed distributions, in comparison to the turbine power and thrust curves. Figure 7 shows the mean wind speed distributions for Europe (including the Baltic and Mediterranean seas), and Venezuela, which was identified above as a country that experiences very high losses. The ambient capacity factors for these regions are very similar (55.8% for Europe, 56.1% for Venezuela), but the CF lost to wakes differs significantly (-5.5% for Europe, -10.1% for Venezuela). This is due to their differing wind speed distributions, as shown in figure 7. Venezuela exhibits a far narrower wind speed distribution than Europe, coinciding with the most wake-sensitive part of the turbine power and thrust curves (figure 8), which is between around 3 and 11 . In Europe, a substantial part of the wind speed distribution sits above 11 , where power output is at its maximum and is insensitive to wake effects.
This influence of the wind speed distribution on wake effects can be more formally explained by considering one-dimensional momentum theory (see e.g. Kulunk (2011)). The power coefficient of a turbine can be related to the axial induction factor , by
Given the wind turbine power curve data for our selected turbine model (see figure 8), can be derived as a function of upstream wind speed. Neglecting wake expansion, the reduction in wind speed downstream of a wind turbine is then given by
@@
Figure 9 shows this estimate as a function of wind speed. This is consistent with our earlier statements regarding the most important wind speed interval for wake effects, between around 3 and 11 . Figure 9 also shows the actual average power lost per turbine by the ‘standard’ farm studied within this work, as a function of wind speed, as simulated using PyWake. Although corresponds to the theoretical power lost by a single turbine downstream of another, neglecting wake expansion, it is clear that it is a good indicator of wake-induced losses at a given wind speed. There appears to be a slight offset in the position of the peak losses, which we attribute to the simplicity of the estimates derived in this section, and in particular the first-order approximation of equation (15), which breaks down around rated speeds where the power curve is highly nonlinear.
To derive a theoretical estimator for the overall mean wake-induced losses, this is combined with the wind speed distribution at a given location , according to
This quantity is very strongly correlated with the simulated wake losses, with an coefficient of 0.96.
It is important to note that this formulation does not depend on the total capacity or layout of the wind farm, but only on the power and thrust curves of the selected turbine. It follows that turbine models with different thrust and power curves will produce a different spatial pattern of wake-induced losses. It may further be possible to mitigate losses by careful selection of turbine types appropriate to the wind speed distribution at the site where they will be deployed. Further exploration of the wake response for different turbine types is beyond the scope of the present study but will be considered in future work.
@@ 3.2 Mapping layout optimisation potential
Figure 11 shows the global distribution of the array layout optimisation potential, which is defined as the percentage improvement in mean power which can be obtained by an optimised layout, compared with the initial square farm layout. There is a negative correlation between wind direction standard deviation and optimisation potential. That is, narrower wind speed distributions are more conducive to array layout optimisation. The greatest optimisation potentials are found in northern South America, particularly on the western (Pacific) coast.
Comparing the optimisation potential of figure 11 with the global distribution of excess wake losses of figure 5, the regions (mostly equatorial) where the wake losses are particularly high also have the greatest optimisation potential. This is also shown in figure 12, which should be compared with figure 3. Mean losses at all ambient CFs are improved by optimisation. Looking at the absolute losses (figs 3(a) and 12(a)), the optimisation has the greatest benefit for ambient CFs around 40–60%. In terms of percentage losses (figs 3(b) and 12(b)), the greatest improvements are found for ambient CFs of around 10–40%.
@@ 3.3 Mapping future changes due to climate
Figure 13(a) shows the mean projected percentage change in ambient capacity factor across the 10 climate models, between the ‘historical’ climate, and 2081–2100 under the SSP585 scenario. Only grid cells with a statistically significant change are shown, where our significance criterion is that at least 8 out of the 10 climate models show the same sign of change. The result is broadly consistent with the global analysis of Ibarra-Berastegui et al. (2023), as well as with other regional studies. Statistically significant decreases in offshore wind potential are observed across both Europe (consistent with (Carvalho et al., 2021)) and the US (consistent with Martinez and Iglesias (2022a)). A decrease in potential is observed offshore of Northern China, which is consistent with Zhang and Li (2021), Deng et al. (2024), although we do not find statistically significant increases in south China. Decreasing potentials are observed on the west coast of India (Basak et al., 2023), as well as most of Australia (Fournier et al., 2023). Our projections around South America are also consistent with the study of de Souza Ferreira et al. (2024).
@@
Figures 14 and 15 show the relationship between the % projected change in ambient CF, and % losses, per grid cell (figure 14), and aggregated per country (figure 15). For the aggregation by country, the significance is determined on a per-country basis, rather than by aggregation over only the significant grid cells belonging to each country. Both indicate a positive correlation between change in ambient CF, and change in % wake losses. This is consistent with the trend found in section 3.1, and in particular figure 3(b).
@@
The majority of grid cells/countries fall into the lower left quadrant of figures 14 and 15, indicating that ambient potential decreases, but that percentage wake-induced losses increase in magnitude, exacerbating the climate change effect. Figure 16 indicates the spatial distribution of these grid cells. These regions include the northern Atlantic, impacting the US, Atlantic Canada, Greenland, Iceland, Ireland and the UK, as well as the US west coast, parts of Australia, India and China, southern Chile, and a number of island states in the southern Pacific. In these regions, decreases in ambient wind potential due to climate change are exacerbated by an increase in the magnitude of the percentage wake-induced losses.
@@
In contrast, countries falling into the upper right quadrant of figures 14 and 15 are projected to both increase in ambient potential and improve in wake losses; these countries include Russia and Canada (driven mostly by strong increases in resource in the Arctic), as well as other isolated regions such as Pakistan and Oman. There are no countries which fall within either the upper left or lower right quadrants for which the statistical significance threshold is met, although there are a small number of individual grid cells meeting these criteria, as shown in figure 14.
@@
Comparing the magnitudes of the various effects studied within this work, the direct impact of climate change on ambient power exceeds magnitudes of 25% for some regions, such as India. The magnitudes of the baseline wake-induced losses (for the current climate) were similar, occasionally exceeding 25% in regions of low ambient potential. Optimisation potentials reached up to 12%, while the indirect climate change impact via wake effects has magnitudes of around 1-2%. However, the relative size and significance of these effects varies by region. For example, there are some regions where optimisation potential exceeds the projected impact of climate change. This serves to further highlight that the challenges facing the offshore wind industry are highly region-dependent.